### v1

@@ 5 Mitigating Anchored Bias
$\mathbf{v}^{9,2859}$ (61.8%) 44.6 48.9 100.0 100.0 86.5 59.4 95.3 95.9 59.4 58.8
$\mathbf{v}^{9,788}$ (51.0%) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 99.4 100.0 97.1 100.0 68.3 25.3 61.9 61.4 21.9 21.3
$\mathbf{v}^{20,1731}(38.3\%)$ 33.6 36.1 80.0 55.6 0.0 0.0 54.3 50.6 20.6 22.5
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 100.0 100.0 100.0 100.0 98.2 100.0 99.5 99.0
$\mathbf{v}^{34,2103}(70.2\%)$ 71.2 75.0 55.0 65.0 94.4 32.3 51.7 59.1 46.8 63.9
$\mathbf{v}^{34,4178}$ (54.7%) 25.1 26.0 8.9 15.0 93.8 35.4 47.4 48.9 49.7 62.9
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 98.2 98.8 100.0 100.0 100.0 100.0 93.9 77.8 95.7 100.0
$\mathbf{v}^{44,2995}$ (44.7%) 24.8 38.8 100.0 100.0 100.0 99.0 96.8 81.0 95.3 96.8
$\mathbf{v}^{44,128}$ (11.9%) 67.8 72.9 96.8 100.0 100.0 100.0 94.9 79.4 94.6 96.8
$\mathbf{v}^{38,4191}$ (100%) 100.0 100.0 100.0 100.0 100.0 94.9 99.6 88.9 97.9 100.0
$\mathbf{v}^{38,4174}$ (20.2%) 11.4 7.1 32.3 33.3 68.5 51.0 22.2 22.2 9.1 12.6
$\mathbf{v}^{37,423}$ (91.8%) 54.3 55.3 100.0 100.0 100.0 92.9 83.0 57.1 41.8 49.5
$\mathbf{v}^{37,2966}$ (67.2%) 34.0 40.0 100.0 100.0 98.4 81.6 79.3 57.1 56.7 56.8
Mitigating anchored bias in attention heads.
[…]
@@
6 Discussion
Is Few-shot learning helpful?
Based on the comprehensive zero-shot learning MCQ across all datasets, we have a good understanding of how important the MLP and attention head are regarding the anchored bias in the GPT2 family. The following question might be whether few-shot learning can mitigate this anchored bias without updating specific value vectors in MLP or recalibrating the attention head. We conduct a preliminary experiment to evaluate GPT2 family across all datasets using 1-shot and 2-shot learning settings. The initial finding is that the anchored bias could be relatively mitigated to a certain extent depending on the length of few-shot prompts and the complexity of the few-shot prompt content. However, the anchored bias still happens, especially for the long and complex few-shot prompts. Even if the anchored bias to ‘A’ disappears, GPT2 model cannot still predict the correct choice and generate other incorrect choices, which needs more investigation. We leave this question as the future work.
Is direct value vector updating in MLP harmful to the ability of GPT2 for other tasks?
We conduct a preliminary experiment to evaluate whether direct value vector updating in MLP is harmful to the ability of GPT2-Small for the original IOI task. The initial result is that modified GPT2-Small can still achieve 85.6% accuracy to successfully predict the non-subject of the last clause. For a better and minimal-harm MLP updating algorithm development, we leave it for future work.
7 Related Work
Mechanistic Interpretability.
Our work falls into the scope of investigating Mechanistic Interpretability, a field studying neural networks with an intent to understand and explain their inner workings. Recent studies have advanced our understanding of how factual associations are managed within transformer models. Geva et al., (2023) investigates the information flow during inference to understand how facts are internally retrieved. Meng et al., 2022a focuses on the storage of knowledge in GPT-like autoregressive transformer models, noting significant differences from masked models, such as unidirectional attention and generation capabilities. They use causal mediation analysis to examine how hidden state activations influence fact recall and introduce a novel model editing method that alters model behaviour, providing deeper insights into the neural mechanisms of information retrieval. Building on these findings, Meng et al., 2022b develops a scalable multi-layer update algorithm that targets transformer weights to enhance memory recall. Similarly, Gupta et al., (2023) proposes a model editing strategy to modify commonsense knowledge in GPT-2 models. Halawi et al., (2023) identifies certain layers of language models following incorrect demonstrations that lead to harmful behaviours. They utilise logit lens that analyses model internal behaviour. Wang et al., (2022) concentrates the task of IOI, and identifies critical circuits of 26 attention heads. Our work also focuses on the task of IOI and reformulates the task as a binary question-answering task to investigate LLM’s positional bias.
Positional Bias of LLMs in MCQ.
Various studies have highlighted the impact of positional bias on the accuracy of LLMs when addressing multiple-choice questions. Pezeshkpour and Hruschka, (2023) demonstrate a significant "sensitivity gap" in LLMs, such as GPT-4, attributing a drop in performance of up to 75% in the zero-shot setting to positional bias. They propose calibration strategies that have successfully improved accuracy by up to 8 percentage points. Similarly, Wang et al., (2023) identify a notable influence of display order on GPT-4’s evaluation scores, advocating a calibration framework that includes Multiple Evidence Calibration, Balanced Position Calibration, and Human In The Loop Calibration. This approach has notably increased evaluation accuracy and better aligned LLM outputs with human judgments. Zheng et al., 2024a explore the "selection bias" of LLMs, where models prefer certain option labels such as "Option A" more frequently. Their novel debiasing technique, PriDe, effectively mitigates this bias by adjusting the prediction distribution at inference time without the need for labels. Wang et al., (2024) further investigate how altering the order of answer options affects LLM understanding, observing significant performance degradation when answer options are reordered. This body of work collectively underscores the necessity for robust strategies to address and correct positional biases in LLMs, ensuring more reliable model performance in multiple-choice settings. Turpin et al., (2024) explores the impact of positional bias of MCQA for Chain-of-thought (CoT) generation. The key findings reveal that the introduction of biasing features significantly affects the CoT predictions. When biasing features are applied, models modify their explanations to support incorrect, bias-consistent answers, occasionally providing sound yet unfaithful reasoning. This suggests need for robust mechanisms to detect and mitigate biases. Zheng et al., 2024b investigate the application of advanced LLMs in the role of evaluators, particularly for assessing responses to open-ended questions. They delve into the strengths and weaknesses of employing LLMs as judges, highlighting biases related to position, verbosity, self-enhancement biases, and constrained logical capabilities. Furthermore, they suggest strategies to alleviate some of these issues.
8 Conclusion
Appendix A The Statistic Information of Each Test Dataset for GPT2 Models
We split each dataset into 90% Infer. set for anchored bias discovering and mitigation, and 10% Eva. set for modified GPT2 model verification on the MCQ task.
Num. GPT2-Small GPT2-Medium GPT2-Large GPT2-XL
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
IOI (2) 410 45 877 97 900 100 773 85
LD (3) 114 12 170 18 180 20 31 3
Greater (4) 289 32 856 95 897 99 883 98
ARC (4) 446 49 748 83 797 88 571 63
CSQA (5) 308 34 720 80 881 97 864 95

### v2

@@ Appendix G Top-10 Unembedded Tokens from Updated Value Vectors
GPT2-Small $\mathbf{v}^{9,788}$ (50.0%) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GPT2-Medium $\mathbf{v}^{20,1731}$ (38.2%) 33.6 33.7 81.8 50.0 0.0 0.0 51.2 51.6 22.4 23.1
GPT2-Large $\mathbf{v}^{34,2103}$ (70.1%) 71.2 63.5 55.6 57.7 32.0 96.2 49.7 50.5 52.9 49.0
$\mathbf{v}^{34,4178}$ (55.4%) 25.1 21.2 12.2 19.2 33.2 93.3 45.2 46.2 55.1 57.7
GPT2-XL $\mathbf{v}^{44,2995}$ (27.5%) 24.8 36.3 100.0 100.0 99.2 100.0 93.3 92.3 92.7 94.2
$\mathbf{v}^{44,128}$ (22.7%) 67.8 64.8 96.8 100.0 100.0 100.0 91.1 90.8 92.5 95.2
$\mathbf{v}^{38,4174}$ (20.2%) 11.4 16.5 38.7 23.1 44.6 75.0 21.0 26.2 7.5 7.7
$\mathbf{v}^{37,423}$ (91.8%) 54.3 60.4 100.0 100.0 93.5 100.0 76.9 81.5 39.5 36.5
$\mathbf{v}^{37,2966}$ (66.2%) 34.0 31.9 100.0 100.0 77.2 97.1 73.7 76.9 55.2 60.6
Table 11: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=7$.
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) 99.5 100.0 100.0 100.0 100.0 100.0 99.8 100.0 100.0 100.0
$\mathbf{v}^{9,2859}$ (61.8%) 22.7 26.9 100.0 100.0 37.8 69.2 89.2 96.2 45.8 41.0
$\mathbf{v}^{9,788}$ (50.0%) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 97.7 95.2 90.6 73.1 17.3 47.1 43.0 52.7 14.2 8.8
$\mathbf{v}^{20,1731}$ (38.2%) 27.6 26.9 61.2 30.8 0.0 0.0 41.4 47.3 13.8 8.8
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 100.0 100.0 100.0 100.0 95.7 95.6 99.2 99.0
$\mathbf{v}^{34,2103}$ (70.1%) 57.6 54.8 40.6 46.2 26.6 85.6 38.6 38.5 42.6 39.4
$\mathbf{v}^{34,4178}$ (55.4%) 16.8 16.3 2.2 3.8 27.0 85.6 35.5 37.4 45.5 45.2
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 97.4 96.7 96.8 92.3 99.8 100.0 87.6 90.8 91.6 92.3
$\mathbf{v}^{44,2995}$ (27.5%) 18.6 30.8 100.0 100.0 97.7 100.0 91.9 92.3 86.9 89.4
$\mathbf{v}^{44,128}$ (22.7%) 59.6 59.3 93.5 92.3 100.0 100.0 90.0 90.8 89.7 92.3
$\mathbf{v}^{38,4191}$ (100%) 99.4 100.0 100.0 100.0 88.0 100.0 96.1 95.4 92.6 89.4
$\mathbf{v}^{38,4174}$ (20.2%) 6.7 7.7 38.7 23.1 37.8 60.6 16.3 18.5 4.7 4.8
$\mathbf{v}^{37,423}$ (91.8%) 38.2 41.8 100.0 100.0 86.9 100.0 67.3 70.8 30.1 29.8
$\mathbf{v}^{37,2966}$ (66.2%) 22.3 26.4 93.5 84.6 71.3 96.2 65.7 64.6 44.6 48.1
Table 12: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=6$.
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) 96.1 96.2 100.0 100.0 94.8 100.0 98.7 98.1 99.4 97.4
$\mathbf{v}^{9,2859}$ (61.8%) 7.6 7.7 92.1 100.0 20.1 38.5 79.8 84.6 25.6 25.6
$\mathbf{v}^{9,788}$ (50.0%) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 92.5 90.4 77.6 26.9 7.8 26.0 29.3 37.4 4.9 2.2
$\mathbf{v}^{20,1731}$ (38.2%) 22.0 20.2 31.2 11.5 0.0 0.0 29.7 36.3 7.1 5.5
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 100.0 100.0 100.0 100.0 93.1 91.2 97.8 97.1
$\mathbf{v}^{34,2103}$ (70.1%) 39.8 37.5 20.6 46.2 14.8 51.9 28.7 29.7 30.8 26.9
$\mathbf{v}^{34,4178}$ (55.4%) 9.8 7.7 0.6 0.0 15.7 51.0 26.1 28.6 35.0 31.7
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 95.2 95.6 93.5 84.6 99.4 100.0 84.2 90.8 85.5 87.5
$\mathbf{v}^{44,2995}$ (27.5%) 12.9 19.8 96.8 92.3 93.1 100.0 87.9 92.3 76.6 85.6
$\mathbf{v}^{44,128}$ (22.7%) 50.7 50.5 90.3 84.6 99.8 100.0 87.6 89.2 84.8 87.5
$\mathbf{v}^{38,4191}$ (100%) 96.9 98.9 100.0 100.0 78.1 100.0 95.1 95.4 82.5 83.7
$\mathbf{v}^{38,4174}$ (20.2%) 3.2 5.5 35.5 23.1 29.9 47.1 13.0 16.9 2.7 1.0
$\mathbf{v}^{37,423}$ (91.8%) 23.2 27.5 100.0 100.0 80.3 100.0 57.8 60.0 20.5 17.3
$\mathbf{v}^{37,2966}$ (66.2%) 12.5 15.4 77.4 76.9 60.7 92.3 56.0 56.9 32.4 29.8
Table 13: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=5$.
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) 76.8 82.7 100.0 100.0 69.2 100.0 93.0 96.2 96.1 94.9
$\mathbf{v}^{9,2859}$ (61.8%) 1.2 3.8 48.2 69.2 4.5 12.8 63.2 69.2 12.3 12.8
$\mathbf{v}^{9,788}$ (50.0%) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 79.8 76.9 50.0 0.0 0.9 1.9 18.9 24.2 0.8 0.0
$\mathbf{v}^{20,1731}$ (38.2%) 15.7 15.4 10.0 0.0 0.0 0.0 20.2 18.7 3.1 2.2
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 100.0 100.0 100.0 100.0 87.6 83.5 95.3 95.2
$\mathbf{v}^{34,2103}$ (70.1%) 23.4 17.3 7.8 26.9 3.6 10.6 20.2 13.2 21.2 14.4
$\mathbf{v}^{34,4178}$ (55.4%) 4.6 2.9 0.0 0.0 5.4 20.2 17.4 16.5 25.1 17.3
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 88.6 90.1 77.4 69.2 95.8 100.0 77.6 84.6 76.0 78.8
$\mathbf{v}^{44,2995}$ (27.5%) 7.2 15.4 90.3 84.6 86.2 98.1 82.7 84.6 61.0 69.2
$\mathbf{v}^{44,128}$ (22.7%) 38.9 41.8 83.9 84.6 98.5 100.0 82.5 86.2 74.8 76.0
$\mathbf{v}^{38,4191}$ (100%) 85.0 87.9 100.0 100.0 68.4 99.0 88.1 87.7 64.4 67.3
$\mathbf{v}^{38,4174}$ (20.2%) 1.3 4.4 32.3 15.4 21.1 35.6 9.3 13.8 1.7 0.0
$\mathbf{v}^{37,423}$ (91.8%) 12.5 17.6 96.8 100.0 69.8 95.2 47.8 47.7 13.5 13.5
$\mathbf{v}^{37,2966}$ (66.2%) 5.6 8.8 64.5 61.5 49.2 73.1 43.4 40.0 21.2 20.2
Table 14: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=4$.
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) 30.7 36.5 58.8 69.2 26.3 59.0 63.9 73.1 64.9 69.2
$\mathbf{v}^{9,2859}$ (61.8%) 0.0 0.0 3.5 0.0 0.0 0.0 35.0 44.2 2.3 2.6
$\mathbf{v}^{9,788}$ (50.0%) 99.5 100.0 100.0 100.0 100.0 100.0 97.8 100.0 99.7 100.0
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 59.4 55.8 16.5 0.0 0.0 0.0 7.6 9.9 0.1 0.0
$\mathbf{v}^{20,1731}$ (38.2%) 11.1 9.6 0.6 0.0 0.0 0.0 10.4 13.2 1.0 0.0
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 100.0 100.0 98.2 100.0 80.7 79.1 86.8 89.4
$\mathbf{v}^{34,2103}$ (70.1%) 10.3 4.8 1.1 0.0 0.1 0.0 10.9 8.8 12.6 5.8
$\mathbf{v}^{34,4178}$ (55.4%) 2.2 1.9 0.0 0.0 0.6 1.0 12.3 9.9 15.9 10.6
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 74.6 76.9 61.3 38.5 85.7 100.0 68.0 76.9 57.4 62.5
$\mathbf{v}^{44,2995}$ (27.5%) 4.1 9.9 77.4 61.5 73.4 96.2 73.9 75.4 42.9 50.0
$\mathbf{v}^{44,128}$ (22.7%) 26.3 31.9 71.0 61.5 92.1 99.0 75.7 78.5 58.9 65.4
$\mathbf{v}^{38,4191}$ (100%) 58.9 57.1 71.0 61.5 58.1 90.4 75.5 75.4 43.2 45.2
$\mathbf{v}^{38,4174}$ (20.2%) 0.5 2.2 32.3 15.4 13.1 22.1 6.8 10.8 1.2 0.0
$\mathbf{v}^{37,423}$ (91.8%) 4.4 3.3 80.6 76.9 54.9 76.0 34.7 38.5 7.5 4.8
$\mathbf{v}^{37,2966}$ (66.2%) 1.9 4.4 51.6 38.5 33.5 54.8 32.2 30.8 11.1 7.7
Table 15: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=3$.
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) 1.7 1.9 0.0 0.0 2.1 10.3 16.8 23.1 14.3 15.4
$\mathbf{v}^{9,2859}$ (61.8%) 0.0 0.0 0.0 0.0 0.0 0.0 8.5 9.6 0.0 0.0
$\mathbf{v}^{9,788}$ (50.0%) 67.6 71.2 71.1 76.9 98.3 100.0 70.6 73.1 92.9 97.4
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 34.7 31.7 0.0 0.0 0.0 0.0 1.5 3.3 0.0 0.0
$\mathbf{v}^{20,1731}$ (38.2%) 8.1 4.8 0.0 0.0 0.0 0.0 4.8 6.6 0.4 0.0
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 95.0 100.0 67.7 100.0 65.1 68.1 70.4 72.1
$\mathbf{v}^{34,2103}(70.1\%)$ 3.1 1.9 0.0 0.0 0.0 0.0 6.1 3.3 5.8 4.8
$\mathbf{v}^{34,4178}$ (55.4%) 0.9 0.0 0.0 0.0 0.0 0.0 6.1 4.4 6.5 5.8
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 51.2 58.2 61.3 38.5 64.2 91.3 52.0 53.8 33.3 34.6
$\mathbf{v}^{44,2995}$ (27.5%) 1.8 4.4 61.3 46.2 49.8 78.8 56.4 58.5 24.0 24.0
$\mathbf{v}^{44,128}$ (22.7%) 12.7 17.6 54.8 53.8 76.6 94.2 61.8 64.6 40.4 42.3
$\mathbf{v}^{38,4191}$ (100%) 21.7 27.5 58.1 38.5 38.1 63.5 54.1 63.1 19.7 21.2
$\mathbf{v}^{38,4174}$ (20.2%) 0.1 0.0 29.0 7.7 6.2 13.5 4.4 6.2 0.8 0.0
$\mathbf{v}^{37,423}$ (91.8%) 1.3 2.2 67.7 53.8 34.7 52.9 22.2 24.6 3.7 3.8
$\mathbf{v}^{37,2966}$ (66.2%) 0.5 1.1 48.4 30.8 19.6 37.5 22.2 18.5 6.6 5.8
Table 16: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=2$.
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) 0.0 0.0 0.0 0.0 0.0 0.0 1.3 1.9 1.0 0.0
$\mathbf{v}^{9,2859}$ (61.8%) 0.0 0.0 0.0 0.0 0.0 0.0 0.9 1.9 0.0 0.0
$\mathbf{v}^{9,788}$ (50.0%) 5.4 7.7 0.0 0.0 24.2 53.8 15.5 19.2 26.9 30.8
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 15.6 16.3 0.0 0.0 0.0 0.0 0.3 0.0 0.0 0.0
$\mathbf{v}^{20,1731}$ (38.2%) 4.2 1.9 0.0 0.0 0.0 0.0 1.7 4.3 0.0 0.0
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 99.9 100.0 66.1 69.2 34.4 100.0 37.5 37.4 39.5 34.6
$\mathbf{v}^{34,2103}$ (70.1%) 0.6 1.0 0.0 0.0 0.0 0.0 2.3 2.2 1.8 1.9
$\mathbf{v}^{34,4178}$ (55.4%) 0.1 0.0 0.0 0.0 0.0 0.0 2.1 1.1 2.7 2.9
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 18.5 17.6 48.4 23.1 33.3 62.5 31.7 33.8 13.8 14.4
$\mathbf{v}^{44,2995}$ (27.5%) 0.6 2.2 48.4 30.8 25.4 49.0 36.1 35.4 9.0 5.8
$\mathbf{v}^{44,128}$ (22.7%) 4.0 7.7 48.4 38.5 48.9 76.9 43.1 38.5 18.9 20.2
$\mathbf{v}^{38,4191}$ (100%) 4.3 6.6 41.9 23.1 16.1 28.8 28.9 29.2 6.5 5.8
$\mathbf{v}^{38,4174}$ (20.2%) 0.0 0.0 22.6 7.7 2.8 4.8 2.5 4.6 0.5 0.0
$\mathbf{v}^{37,423}$ (91.8%) 0.4 1.1 48.4 30.8 15.6 32.7 12.3 12.3 1.5 1.9
$\mathbf{v}^{37,2966}$ (66.2%) 0.1 0.0 32.3 15.4 7.5 15.4 11.2 15.4 2.0 0.0
Appendix I Top-1 Classification Accuracy Changes Using Attention Pattern Recalibration
We also evaluate the top-1 classification accuracy when the attention pattern recalibration is used to the identified attention head. In Table 17, we can find that most top-1 classification accuracy is close to 0%, which indicates that those identified attention heads play less important roles compared to specific MLP modules in each GPT2 model under the MCQ task setting.
@@ Appendix I Top-1 Classification Accuracy Changes Using Attention Pattern Recalibration
Appendix J One-shot and Two-shot MCQ Classification Results
Based on the discussion in § 6, we further conduct an experiment to evaluate whether the anchored bias can be mitigated under the few-shot learning setting. Fig. 7 and Fig. 8 show that 2-shot leaning performs better than 1-shot setting across all datasets, especially for IOI dataset. However, GPT2-XL always struggle to predict higher accuracy across all datasets. This finding indicates that simple few-shot learning cannot mitigate anchored bias and a more comprehensive analysis is needed to investigate this bias.
Figure 7: One-shot MCQ classification accuracy of different datasets across GPT2 family.
Figure 8: Two-shot MCQ classification accuracy of different datasets across GPT2 family.
Appendix K Damage to Updated GPT2 Family’s Performance on IOI and Greater Datasets
Based on the discussion in § 6, we further evaluate the damage of direct value vector update from MLP for the general ability of GPT2 family. We choose original IOI and Greater-than datasets as they have been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024; Hanna et al.,, 2024). Fig. 9 and Fig. 10 show that classification accuracy is acceptable when a value vector with a specific layer and dimensionality is updated, which matches the findings from Gu et al., (2024).
Figure 9: Damage to updated GPT2 family’s classification accuracy on original IOI dataset using Eq. 8 with $\lambda_{2}=8$.
Figure 10: Damage to updated GPT2 family’s classification accuracy on original Greater dataset using Eq. 8 with $\lambda_{2}=8$.
Appendix L MCQ Prompt Template with Random Words
In order to evaluate whether anchored bias is sensitive to the specific content of the input MCQ prompts, we built two random MCQ datasets. One of them contains randomly selected characters and we concatenate them as a word with length from 5 to 10, and each MCQ prompt is built based on “Question: <Question sample> Answer Choices: <Multiple Choices> Answer:”. For the random word dataset, we randomly select words from a word vocabulary [^9] https://www.mit.edu/~ecprice/wordlist.10000. Finally, each dataset has 80 samples with the number of choices from 2 to 5 (see Table 19). As shown in Table 18, anchored bias still happens across GPT2 family, especially for GPT2-Large and GPT2-XL. This finding confirms our findings that GPT2 family exhibit the anchored bias with significant regularity across various MCQ datasets.
Table 18: The percentage of anchored bias occurring when MCQ prompt template is replaced with random characters and random words.
Type GPT2-Small GPT2-Medium GPT2-Large GPT2-XL
Random Characters 70.0 69.0 99.0 96.0
Random Words 65.0 82.0 95.0 100.0
Table 19: The MCQ prompt template with random characters and random words.
Random Characters Question: cgryp rkodmajc ajwsqby jnqqaqcmsa agbyyasj ibngbxtn dazvqigre urbnmumw ltpjslayp ighfudgy hbaldde? Answer Choices: A: samdaheepw ltvlpeh B: rqqtxdgiyb rznosxhk djpsitdar ubjgq ioamje C: bkdjziiy Answer:
Random Words Question: citysearch logical bidder discount kentucky forming rapid digit flash putting reid liechtenstein mate? Answer Choices: A: seven owner voluntary B: clicking it C: harassment beam firewire D: run helpful Answer:
Appendix M Full Circuit of Anchored Bias of Each GPT2 Model
Based on the identified MLP module and attention head using logit difference between anchored choice ‘A’ and correct choices ‘B,C,D,E’. We can construct the full anchored-bias circuit for each GPT2 model. Fig. 11,12,13 show the full anchored-bias circuit of GPT2-Medium, GPT2-Large and GPT2-XL including identified attention heads and MLPs starting from layer 0 to the final layer.
Figure 11: The full circuit of anchored bias for GPT2-Medium model, where each attention head and MLP module are selected when MLP and attention pattern logit difference threshold is larger than 4. The percentage within each module indicates the probability of anchored bias across different datasets for GPT2-Medium model when the threshold is larger than 4.
Figure 12: The full circuit of anchored bias for GPT2-Large model, where each attention head and MLP module are selected when MLP and attention pattern logit difference threshold is larger than 4. The percentage within each module indicates the probability of anchored bias across different datasets for GPT2-Large model when the threshold is larger than 4.
Figure 13: The full circuit of anchored bias for GPT2-XL model, where each attention head and MLP module are selected when MLP and attention pattern logit difference threshold is larger than 4. The percentage within each module indicates the probability of anchored bias across different datasets for GPT2-XL model when the threshold is larger than 4.